In the month of March, Blossom Salon services 550 clients at an average price of $100. During the month, fixed costs were $17,520 and variable costs were 60% of sales.

Determine the total contribution margin in dollars, the per unit contribution margin, and the contribution margin ratio.

Contribution margin in dollars

$

Contribution margin per unit

$

Contribution margin ratio %

Using the contribution margin technique, compute the break-even point in dollars and in units.

Break-even sales

$

Break-even sales units
Answer #1
Answer
Contribution margin in dollars (WN-1) $            22,000
Contribution margin per unit (WN-2) $ 40 Per units
Contribution margin ratio% ( WN-3) 40%

Break-even sales ( WN-4)

$            43,800
Break-even sales Units (WN-5) $ 438    Unit
Working Note
WN- 1 Contribution Margin = Contribution Margin Per Unit * Total Sales Volume
$40*550 = $22,000
WN-2

Contribution Margin Per Unit= Selling Price Per Unit - Variable Cost Per Unit

100 -( 0.60*100) = $40 per Units
WN-3 Contribution Margin Ratio = Contribution Margin / Sales *100
40/100*100 = 40%
WN-4 Break Even Sales (in Dollars) = Fixed Cost / Contribution Margin Ratio
17520/40%=$43,800
Wn-5 Break Even Sales (in Units) = Fixed Cost / Contribution Margin Per Unit
17520/40= 438 Units

> this is correct dont be scared becasue of the thumbs down its correct

Lil_lostcar Sun, Apr 9, 2023 3:18 PM